OpenStudy (anonymous):

3 x what = 13??? Use decimals if needed

OpenStudy (gorv):

let that is x

OpenStudy (gorv):

3*x=13

OpenStudy (gorv):

divide both side by 3

OpenStudy (gorv):

what u got ??

OpenStudy (anonymous):

I can't use x for it because it is asking for a number, not a letter..

OpenStudy (gorv):

loll..we will find the value of x ..and the value of x will be the number we need

OpenStudy (anonymous):

you will get a number in the end...but first use a letter

OpenStudy (confluxepic):

First divide 13 by 3.

OpenStudy (anonymous):

Okay. 13/3 = 4.3?

OpenStudy (confluxepic):

Yeah. That is the value of the number.

OpenStudy (gorv):

well done @MMV04

OpenStudy (anonymous):

Thanks. :)

OpenStudy (anonymous):

3x = 13 --- divide both sides by 3...because what you do on one side of the equal sign, you have to do on the other side 3x/3 = 13/3 (the 3's on the left side cancel out) x = 13/3 x = 4.3 there you go...step by step you can even check your answer by subbing in 4.3 for x 3(4.3) = 13 12.9 (rounds to 13) = 13 (correct)